OpenStudy (anonymous):

The owner of a bike shop sells unicycles and bicycles and keeps inventory by counting seats and wheels. One day, she counts 20 seats and 28 wheels. Which equation in slope-intercept form shows that x bicycles and y unicycles together have a total of 28 wheels

OpenStudy (anonymous):

A. Y=-2x+28 B. Y=-2x-28 C. Y=2x-28 D. Y=2x+28